OpenStudy (anonymous):

x-y=0 -3x-y=2 Use linear combinations to solve the system of linear equations.

OpenStudy (anonymous):

couldn't you just use substitution method to solve the linear equation? x-y=0 -3x-y=2 x=y -3(y)-y=2 -4y=2 y=(-1/2)

OpenStudy (anonymous):

thank you so much can you help me with one more problem v-w=-5 v+2w=4